# Learning from core-collapse supernova remnants on the explosion mechanism

### v1

@@ 3 The required numerical resolution
In a powerful magnetic activity $v_{a}\simeq v_{\rm Kep}$, the Alfven speed will not be larger than the rotation speed. Considering then this ratio of $v_{a}/v_{\rm Kep}$ and that the reconnection zone should be resolved by more than two cells, better at least four cells, I conclude that for simulating the JJEM, the grid cells should be $\Delta\lesssim 0.002r$.
In the recent magnetohydrodynamical simulations of CCSNe that Varma et al. (2023), Powell & Müller (2024), and Zha et al. (2024b), conducted, the numbers of grid cells in the three coordinates were $(N_{r},N_{\theta},N_{\phi})=(550,128,256)$, corresponding to an angular resolution of $1.4^{\circ}$, or $\Delta_{\theta,\phi}=0.0245r$. They have near the NS $\Delta_{r}\simeq 0.015r$ in the radial direction. This resolution is several times, and up to an order of magnitude, too coarse for the demand of the JJEM. Raynaud et al. (2020) performed magnetohydrodynamic simulations of magnetar formation. They have high resolution but simulated only the inner $\simeq 40{~{}\rm km}$, namely, the proto-NS but not its surroundings. With the grid of $(N_{r},N_{\theta},N_{\phi})=(257,512,1024)$, the resolution of $\Delta_{\theta,\phi}=0.006r$ is almost as required for simulating the JJEM, but not yet as needed.
@@ 4 Discussion and Summary
Last year, studies of the JJEM analyzed in depth the point-symmetrical morphologies of about ten CCSNRs (see list in Section 1). The main conclusion of the in-depth analysis is that jittering jets shaped all these CCSNRs, and the shaping jets were part of the tens of jets that exploded the star (e.g., Soker & Shishkin 2024). The conclusion of jet-shaping of CCSNRs is based in part on carefully studying many morphological similarities between CCSNRs and jet-shaped planetary nebulae and hot gas in clusters of galaxies where pairs of jets shape pairs of bubbles, clumps, rims, and nozzles (e.g., Soker 2024b; Bear et al. 2024; see a talk from 2024: https://www.youtube.com/watch?v=hJYc6EfgxJU). Although the tool of identifying the morphological imprints of jets is not familiar in the CCSN community, it is standard in analyzing structures of planetary nebulae and hot gas in clusters of galaxies, making the claim for jet-shaped CCSNRs robust. The analysis of these point-symmetrical CCSNRs motivated the present paper.
@@
Based on the solidification of the JJEM in 2024 and the understanding that jets shape many CCSNRs, I estimated more parameters of the JJEM (rows 12-14 in Table 1). I estimated the typical half opening of the shaping jittering jets to be $\alpha_{\rm j}\simeq 1^{\circ}-10^{\circ}$. Namely, these are not wide jets. I also estimated that an accretion disk around the newly born NS can launch jets even when its lifetime is only several times the orbital period in the radius where it launches the jets. Table 1 lists the estimated parameters of the JJEM.
@@
While studies of the JJEM in 2024 focused on point-symmetrical morphologies of CCSNRs, studies of the competing delayed-neutrino explosion mechanism have focused on three-dimensional simulations to revive the stalled shock (e.g., Burrows et al. 2024a; Janka & Kresse 2024; Müller 2024; Müller et al. 2024; Nakamura et al. 2024; van Baal et al. 2024; Wang & Burrows 2024). As far as I know, when I wrote this paper, none of the studies of the delayed-neutrino explosion mechanism suggested any explanation for point-symmetrical CCSNRs. Some researchers in private communications mentioned shaping by post-explosion jets and/or interaction with the interstellar medium or the circumstellar material. However, in a detailed analysis of several CCSNRs, Soker & Shishkin (2024) find that these mechanisms encounter severe difficulties. Post-explosion jets and ambient gas cannot explain the point-symmetrical structures of some heavy elements that come from the deep core of the CCSN progenitor (e.g., O, Ne, Mg, Si, S, Fe) and as observed in some CCSNRs, and nor the point-symmetrical structures in the inner zones of some CCSNRs.
@@
The strongest argument against the JJEM made by supporters of the delayed neutrino explosion mechanism is that their simulations do not obtain jittering jets. In this study, I dismissed this argument. In Section 3, I argued that the numerical resolution in magnetohydrodynamical simulations appropriate to the JJEM should be much higher than in current simulations. Specifically, the typical size of grid cells in radius $r$ should be $\Delta\lesssim 0.002r$. Current CCSN simulations have grid cells larger by $\gtrsim 3-5$ times from this constraint. I conclude that the simulations of the CCSN explosion mechanism are still a long way from simulating the correct CCSN explosion mechanism.
@@ Acknowledgements
A grant from the Pazy Foundation supported this research.

### v2

@@ 4 The required numerical resolution
In a powerful magnetic activity $v_{a}\simeq v_{\rm Kep}$, the Alfven speed will not be larger than the rotation speed. Considering then this ratio of $v_{a}/v_{\rm Kep}$ and that the reconnection zone should be resolved by more than two cells, at least five cells (e.g., Figueiredo et al. 2024; Morillo & Alexakis 2025), I conclude that for simulating the JJEM, the grid cells should be $\Delta\lesssim 0.002r$.
The constraint on the required resolution for treating reconnection seems to be stronger than the above demand of $\Delta\lesssim 0.05\lambda_{\rm tube}$. In a very recent study of magnetohydrodynamical simulations of the reconnection process, Morillo & Alexakis (2025) conducted simulations with resolutions of $\Delta\simeq 0.01\lambda_{\rm tube}$ to $\Delta\simeq 0.0003\lambda_{\rm tube}$. They conclude that only the very high resolution, $\Delta<0.001\lambda_{\rm tube}$ for the parameters they present (e.g., their figure 5), start to converge. Moreover, some studies consider the ion inertial length, $d_{i}=\sqrt{m_{i}c^{2}/4\pi n_{i}e^{2}}$, to be the relevant scale to resolve (e.g., Shay et al. 2024; Liu et al. 2025 for recent reviews). Taking protons to be the dominant ions, with a mass $m_{i}=m_{p}$ and a charge $q=e$, and scaling with a proton number density of $n_{i}$ corresponding to about $\rho\simeq 1{~{}\rm g}{~{}\rm cm}^{-3}$, yields $d_{i}\simeq 2\times 10^{-5}(n_{i}/10^{24}{~{}\rm cm}^{-3})^{-1/2}{~{}\rm cm}$. For the solar wind at earth $n_{i}\approx 1{~{}\rm cm}^{-3}$ and $d_{i}({\rm Earth)}\approx 100{~{}\rm km}\approx 0.01R_{\earth}$, while in the solar corona $n_{i}\approx 10^{9}{~{}\rm cm}^{-3}$ and $d_{i}({\rm Coro})\approx{\rm few~{}meters}\approx 10^{-8}R_{\odot}$. The density behind the stalled shock at about $\simeq 150{~{}\rm km}$ is $\rho\gtrsim 10^{8}{~{}\rm g}{~{}\rm cm}^{-3}$ (e.g., Janka 2001), yielding $d_{i}({\rm SN})\lesssim 10^{-8}{~{}\rm cm}\approx 10^{-14}R_{\rm NS}$. Overall, the condition on the minimum resolution of $\Delta\lesssim 0.002r$ might be a very optimistic one, and it might be that the required resolution is extremely demanding, implying the need for subgrid physics, namely, introducing reconnection by hand.
In the recent magnetohydrodynamical simulations of CCSNe that Varma et al. (2023), Powell & Müller (2024), and Zha et al. (2024b), conducted, the numbers of grid cells in the three coordinates were $(N_{r},N_{\theta},N_{\phi})=(550,128,256)$, corresponding to an angular resolution of $1.4^{\circ}$, or $\Delta_{\theta,\phi}=0.0245r$. They have near the NS $\Delta_{r}\simeq 0.015r$ in the radial direction. This resolution is several times, and up to an order of magnitude, too coarse for the demand of the JJEM. The same holds for other recent simulations of CCSNe; e.g., Wang & Burrows (2024) have a grid of $(N_{r},N_{\theta},N_{\phi})=(10240,128,256)$, and Janka & Kresse (2024), whose best angular resolution is $2^{\circ}$. Raynaud et al. (2020) performed magnetohydrodynamic simulations of magnetar formation. They have high resolution but simulated only the inner $\simeq 40{~{}\rm km}$, namely, the proto-NS but not its surroundings. With the grid of $(N_{r},N_{\theta},N_{\phi})=(257,512,1024)$, the resolution of $\Delta_{\theta,\phi}=0.006r$ is almost as required for simulating the JJEM, but not yet as needed.
@@ 5 Discussion and Summary
Over the last two years, studies of the JJEM have analyzed in depth the point-symmetrical morphologies of fifteen CCSNRs (see the list in Section 1). The main conclusion of the in-depth analysis is that jittering jets shaped all these CCSNRs, and the shaping jets were part of the tens of jets that exploded the star (e.g., Soker & Shishkin 2025). The conclusion of jet-shaping of CCSNRs is based in part on carefully studying many morphological similarities between CCSNRs and jet-shaped planetary nebulae and hot gas in clusters of galaxies where pairs of jets shape pairs of bubbles, clumps, rims, and nozzles (e.g., Soker 2024b; Bear et al. 2025; see a talk from 2024: https://www.youtube.com/watch?v=hJYc6EfgxJU, and 2025: Soker 2025c). Although the tool for identifying the morphological imprints of jets is not familiar in the CCSN community, it is standard in analyzing structures of planetary nebulae and hot gas in clusters of galaxies, making the claim for jet-shaped CCSNRs robust. The analysis of these point-symmetrical CCSNRs motivated the present paper.
@@
I estimate some typical properties of the JJEM to distinguish it from competing supernova explosion models (rows 12-14 in Table 1). I estimated the typical half opening of the shaping jittering jets to be $\alpha_{\rm j}\simeq 1^{\circ}-10^{\circ}$. Namely, these are not wide jets. I also estimated that an accretion disk around the newly born NS can launch jets even when its lifetime is only several times the orbital period in the radius where it launches the jets. Table 1 lists the estimated parameters of the JJEM.
@@
While studies of the JJEM in 2024-2025 focused on point-symmetrical morphologies of CCSNRs, studies of the competing delayed-neutrino explosion mechanism have focused on three-dimensional simulations to revive the stalled shock (e.g., Burrows et al. 2024a; Janka & Kresse 2024; Müller 2024; Müller et al. 2024; van Baal et al. 2024; Wang & Burrows 2024; Nakamura et al. 2025). As far as I know, when I wrote this paper, none of the studies of the delayed-neutrino explosion mechanism suggested any explanation for point-symmetrical CCSNRs. Some researchers in private communications mentioned shaping by post-explosion jets and/or interaction with the interstellar medium or the circumstellar material. However, in a detailed analysis of several CCSNRs, Soker & Shishkin (2025) find that these mechanisms encounter severe difficulties. Post-explosion jets and ambient gas cannot explain the point-symmetrical structures of some heavy elements that come from the deep core of the CCSN progenitor (e.g., O, Ne, Mg, Si, S, Fe) and as observed in some CCSNRs, and nor the point-symmetrical structures in the inner zones of some CCSNRs.
Section 3 compared the two explosion mechanisms I studied to some observational properties of CCSNe. Table 2 summarizes the main observational properties. I find that the JJEM does much better than the neutrino-driven mechanism in accounting for observations.
@@
The strongest argument against the JJEM made by supporters of the delayed neutrino explosion mechanism is that their simulations do not obtain jittering jets. In this study, I dismissed this argument. In Section 4, I argued that the numerical resolution in magnetohydrodynamical simulations appropriate to the JJEM should be much higher than in current simulations. Specifically, the typical size of grid cells in radius $r$ should be $\Delta\lesssim 0.002r$. Current CCSN simulations have grid cells that are larger by $\gtrsim 3-5$ times the constraint. I conclude that the simulations of the CCSN explosion mechanism still have a long way to go before accurately simulating the correct CCSN explosion mechanism.
@@ Acknowledgements
I thank an anonymous referee for useful comments that added significantly to the paper. A grant from the Pazy Foundation supported this research. I thank the Charles Wolfson Academic Chair at the Technion for the support.